# Adaptive Quantum Control

### v1

@@ Abstract
This paper presents a model-free, learning-based adaptive controller for the density tracking problem in a two-level Lindblad-Gorini-Kossakowski-Sudarshan (LGKS) quantum system. The adaptive controller is based on the continuous-time retrospective cost adaptive control. To preserve the geometric properties of the quantum system, an adaptive PID controller driven and optimized by Ulhmann’s fidelity is used. The proposed controller is validated in simulation for a low and a high-entropy density-tracking problem.
keywords: adaptive control, quantum control, density-tracking.
@@ I INTRODUCTION
Quantum dynamical systems have several applications in sensing, metrology, communication, computation, and the internet [giovannetti2011advances, toth2014quantum, taylor2016quantum, cacciapuoti2020entanglement, nielsen2001quantum]. The reliable operation of quantum circuits, such as quantum logic gates, requires the precise and predictable behavior of quantum dynamical systems, much like how classical digital circuits depend on the accurate design of their fundamental electrical components, such as resistors and transistors [barenco1995elementary]. In quantum computing, the basic unit of information is the qubit, which, unlike classical bits limited to two distinct states (00 or $1$), can exist in a superposition of both states, enabling quantum algorithms to perform computations exponentially faster than classical systems [grover1996fast, shor1999polynomial]. The successful realization of quantum devices depends on the precise control of quantum dynamical systems.
Quantum control involves manipulating quantum systems to achieve specific outcomes. It has applications in fields such as quantum computing, communication, and sensing. This is achieved through external controls, such as electromagnetic fields, to direct the evolution of quantum systems. The nature of the desired quantum event in a control problem depends on the quantum problem objective. For instance, in applications such as quantum computing, metrology, communication, and optical systems [yamamoto2024engineering, zaletel2021preparation], a low entropy state is desired to reduce the randomness and uncertainties from the output, whereas high entropy is desired when high randomness is needed in quantum phenomena. Applications such as heat spin engines, quantum cryptography, and quantum thermodynamics simulation [von2019spin, Sehgal2021, somhorst2023quantum], can be conceived as quantum control problems of this category. Quantum control strategies often rely on optimizing the system’s time evolution, governed by quantum operators such as Hamiltonians and density matrices. The Schrödinger equation is used for closed systems, while open systems interacting with the environment follow the LGKS equation [brif2010control]. Recent advancements have improved control techniques for both isolated and open quantum systems [dong2010quantum, cong2014control]. However, due to fundamental physical limitations of quantum state measurement and uncertainty in the knowledge of quantum dynamics, the performance achievable by model-based control systems is limited.
Adaptive quantum control addresses the challenges posed by fundamental physical limitations in quantum state measurement and uncertainties in quantum dynamics. This control strategy adjusts its parameters based on feedback from the quantum system, compensating for variations in system parameters and external disturbances[Kosut2003, kosut2013adaptive, Liu2020]. By incorporating online learning algorithms, adaptive control reduces the impact of incomplete system knowledge and measurement limitations [egger2014adaptive] This approach is significant for practical quantum technologies, including quantum error correction and quantum information processing, where uncertainties affect performance.
Despite these advancements, existing adaptive controllers primarily address parameter uncertainty while still requiring some knowledge of the quantum system model, which can be difficult to obtain in complex quantum systems. This fundamental assumption enables adaptive controllers in the literature to focus mainly on managing parameter uncertainties [egger2014adaptive, Koswara2014, Hu2016].
This paper considers the problem of designing an adaptive controller for an open quantum system without any prior knowledge of the system’s dynamics. In this preliminary work, we consider Uhlmann’s fidelity as the error metrics [UHLMANN1976273]. Uhlmann’s fidelity is a unique real scalar metric that allows one to compare the similarity of two quantum states. In this paper, we design a continuous-time adaptive controller based on the Retrospective Cost Adaptive Control (RCAC) Technique [ali2015retrospective, Goel2015, Sobolic2016, poudel2023learning] for the tracking problem of an open quantum system whose dynamics are governed by the Lindblad-Gorini-Kossakowski-Sudarshan (LGKS) equation [manzano2020short]. The contributions of this paper are thus 1) the design of a continuous-time adaptive controller optimized by RCAC for the open quantum system described by the LGKS equation without prior knowledge of the system’s dynamics, 2) the design of an adaptive PID controller driven by the Uhlmann’s fidelity, and 3) numerical validation of the proposed controller in a low and high entropy state-tracking problem.
@@
The paper is organized as follows. Section II describes the LGKS equation in the context of a controlled quantum open system. Section III presents the mathematical foundations for continuous time RCAC with forgetting factors. Section IV describes the numerical simulation that validate the application of the adaptive controller in a two-level open system. Finally, the paper concludes with a discussion of results and future research directions in section V.
@@ II Problem Formulation
Consider the Lindblad-Gorini-Kossakowski-Sudarshan equation [manzano2020short]
$\dot{\rho}$ $=-i\hbar[H,\rho]+\sum_{i=1}^{n}\left(L_{i}\rho L_{i}^{\rm H}-\dfrac{1}{2}\{L_{i}^{\rm H}L_{i},\rho\}\right),$ (1)
where $\rho\in\mathbb{C}^{2\times 2}$ is the density matrix, the system Hamiltonian $H\in\mathbb{C}^{2\times 2}$ is Hermitian, $n$ is the number of damping terms, and for $i\in(1,2,\ldots,n),$ the matrices $L_{i}\in{\mathbb{C}}^{2\times 2}$ are the jump operators. Without loss of generality, the units are chosen such that $\hbar=1.$ The diagonal elements of $\rho$ represent the probability of a quantum event. Thus, in the case where $\rho\in{\mathbb{C}}^{2\times 2},$ (1) is called a two-level quantum dynamic system.
It is shown in [breuer2002theory, p 123] that, if ${\rm tr}\,\rho(0)=1,$ then, for all $t>0,$ $\rho(t)$ given by (1) satisfies ${\rm tr}\,(\rho(t))=1.$ Furthermore, for all $t\geq 0,$ $\rho(t)$ is positive semidefinite and Hermitian.
As shown in [altafini2012modeling], the Hamilotonian $H$ can be decomposed as
@@
where $H_{0}\in{\mathbb{C}}^{2\times 2}$ is the free Hamiltonian, $H_{1}\in{\mathbb{C}}^{2\times 2}$ is the control Hamiltonian, and $u$ is the control signal to be designed. In this paper, we consider the case where $u\in{\mathbb{R}},$ which ensures that the Hamiltonian $H$ is Hermitian at all time instants.
The control objective is to design an adaptive feedback control law such that the density matrix $\rho$ converges to the desired density matrix $\rho_{\rm d}.$
@@ II-A Density Error
As shown in [jones2024controlling], the Uhlmann–Jozsa fidelity between two density matrices $\rho_{1},\rho_{2}$ is defined as
@@
It follows from Theorem A.5 that $F(\rho_{1},\rho_{2})\in[0,1].$ Furthermore, $F(\rho_{1},\rho_{2})=1$ if and only if $\rho_{1}=\rho_{2}.$ Finally, as shown in [jozsa1994fidelity], in the case where $\rho_{1},\rho_{2}\in\mathbb{C}^{2\times 2},$ the Uhlmann–Jozsa fidelity is given by
@@
In this paper, we consider the density error
@@
where $\rho_{\rm{d}}\in\mathbb{C}^{2\times 2}$ is the desired density
@@ II-B Adaptive PID Control
An adaptive proportional-integral-derivative controller can be written as
@@
where $k_{\rm p}(t),k_{\rm i}(t),$ and $k_{\rm d}(t)$ are the adaptive proportional, integral, and derivative gains. Note that the controller can be written as
@@
$\Phi(t)\stackrel{{\scriptstyle\triangle}}{{=}}\begin{bmatrix}e(t)&\gamma(t)&\dot{e}(t)\end{bmatrix},\quad\theta(t)\stackrel{{\scriptstyle\triangle}}{{=}}\begin{bmatrix}k_{\rm p}(t)\\ k_{\rm i}(t)\\ k_{\rm d}(t)\end{bmatrix}.$ (8)
Note that integral state $\gamma(t)$ satisfies $\dot{\gamma}(t)=e(t).$
The objective of the adaptive controller is to update the controller gains and thus compute the control signal $u(t)$ such that the density error $e(t)$ converges to zero. In this paper, we use the continuous-time retrospective cost adaptive control algorithm, described in the next section, to update the controller gains.
@@ III Continuous Time Retrospective Cost Adaptive Control
where $x(t)\in\mathbb{C}^{l_{x}}$ is the state, $u(t)\in\mathbb{C}^{l_{u}}$ is the input, $y(t)\in\mathbb{C}^{l_{y}}$ is the measured output, and the vector functions $f$ and $g$ are the dynamics and the output maps. Define the performance variable
@@
where the regressor matrix $\Phi(t)\in\mathbb{C}^{l_{u}\times l_{\theta}}$ contains the measured data and the vector $\theta(t)\in\mathbb{C}^{l_{\theta}}$ contains the controller gains to be optimized. Various linear parameterizations of MIMO controllers are described in [goel_2020_sparse_para].
@@
$\Phi_{\rm f}(t)$ $\stackrel{{\scriptstyle\triangle}}{{=}}G_{\rm f}(s)\left[\Phi(t)\right]\in\mathbb{C}^{l_{y}\times l_{\theta}},$ (14)
$u_{\rm f}(t)$ $\stackrel{{\scriptstyle\triangle}}{{=}}G_{\rm f}(s)\left[u(t)\right]\in\mathbb{C}^{l_{y}\times l_{\theta}},$ (15)
@@
$J(t,\hat{\theta})$ $=\int_{0}^{t}\Big{(}e^{-\lambda(t-\tau)}\hat{z}(\tau)^{\rm T}R_{z}\hat{z}(\tau)$
$\quad+(\Phi(\tau)\hat{\theta})^{\rm T}R_{u}(\Phi(\tau)\hat{\theta})\Big{)}{\rm d}\tau$
$\quad+e^{-\lambda t}\hat{\theta}^{T}R_{\theta}\hat{\theta},$ (20)
where $R_{z},R_{u},$ and $R_{\theta}$ are positive definite weighting matrices of appropriate dimensions, and $\lambda<0$ is the exponential forgetting factor.
@@
Consider the cost function $J(t,\hat{\theta})$ given by (20). For all $t\geq 0,$ define the minimizer of $J(t,\hat{\theta})$ by
$\theta(t)\stackrel{{\scriptstyle\triangle}}{{=}}\underset{\hat{\theta}\in{\mathbb{R}}^{l_{\theta}}}{\operatorname{argmin}}J(t,\hat{\theta}).$ (21)
@@
Next, define, for all $\geq 0,$ $P(t)\stackrel{{\scriptstyle\triangle}}{{=}}A(t)^{-1}.$ Using the fact that $P(t)A(t)=I,$ it follows that

### v2

# Adaptive Fidelity-Based Density Tracking for Open Quantum Systems

@@ Abstract
This paper presents an online learning-based adaptive control framework for density-matrix tracking in a two-level Lindblad–Gorini–Kossakowski–Sudarshan (LGKS) quantum system, in which the feedback control law does not require prior knowledge of the system Hamiltonian or dissipative operators. The adaptive controller is based on a continuous-time formulation of retrospective cost adaptive control (RCAC). To preserve the geometric structure of the quantum-state evolution, an adaptive PID controller driven by Uhlmann’s fidelity is employed. The proposed approach is validated in numerical simulations for both low-entropy and high-entropy density-tracking tasks, and robustness to measurement noise in the feedback path is investigated.
keywords: Quantum feedback control, open quantum systems, online learning, adaptive PID control, Lindblad dynamics
@@ I INTRODUCTION
Quantum dynamical systems have several applications in sensing, metrology, communication, computation, and the internet [10, 30, 29, 6, 23]. The reliable operation of quantum circuits, such as quantum logic gates, requires the precise and predictable behavior of quantum dynamical systems, much like how classical digital circuits depend on the accurate design of their fundamental electrical components, such as resistors and transistors [3]. In quantum computing, the basic unit of information is the qubit, which, unlike classical bits limited to two distinct states (0 or $1$), can exist in a superposition of both states, enabling quantum algorithms to perform computations exponentially faster than classical systems [12, 26]. The successful realization of quantum devices thus depends on the precise control of quantum dynamical systems.
Quantum control concerns the manipulation of quantum systems to achieve specific dynamical or state-transfer objectives and plays a central role in quantum computing, communication, and sensing. Such control is typically realized through externally applied fields, for example, electromagnetic control inputs, that steer the system’s evolution. The desired quantum state depends on the application. In quantum computing, metrology, communication, and optical systems [35, 36], low-entropy (high-purity) states are often desirable to preserve coherence and reduce uncertainty, whereas in other settings, high-entropy or mixed states may be of interest when intrinsic randomness or thermal effects are central to the application. Examples of the latter include heat engines based on spin systems, quantum cryptography, and quantum thermodynamics simulations [32, 25, 28]. Quantum control strategies typically seek to shape the system’s time evolution, governed by quantum operators such as Hamiltonians and density matrices. Closed quantum systems evolve according to the Schrödinger equation, whereas open quantum systems interacting with an environment are commonly modeled using the Lindblad–Gorini–Kossakowski–Sudarshan (LGKS) equation [5]. Recent years have seen significant progress in control techniques for both isolated and open quantum systems [8, 7].
In practice, the performance of model-based quantum control strategies is limited by uncertainty in the underlying system dynamics, unmodeled environmental effects, and constraints on the accuracy and availability of quantum state measurements. These challenges motivate the development of adaptive quantum control methods that adjust controller parameters online based on system feedback to compensate for modeling errors, parameter drift, and disturbances [17, 18, 20]. By incorporating online learning mechanisms, adaptive control can mitigate the impact of incomplete system knowledge and imperfect measurements [9]. Such approaches are relevant for practical quantum technologies, including quantum error correction and quantum information processing, where uncertainties and environmental interactions directly affect performance.
Despite these advances, most existing adaptive quantum control approaches primarily address parametric uncertainty while still requiring partial knowledge of the system model, such as the Hamiltonian structure or dissipative operators, which may be difficult to obtain accurately for complex quantum systems. Consequently, much of the existing literature focuses on robustness to parameter variations rather than control design in the absence of an explicit dynamical model [9, 19, 14]. This motivates the investigation of control architectures in which the feedback law itself does not rely on detailed knowledge of the system dynamics.
This paper considers the problem of designing an adaptive controller for an open quantum system in which the control law does not require prior knowledge of the system Hamiltonian or dissipative operators. Specifically, we develop a continuous-time adaptive controller based on the Retrospective Cost Adaptive Control (RCAC) framework [1, 11, 27, 24] for density-matrix tracking in open quantum systems governed by the Lindblad–Gorini–Kossakowski–Sudarshan (LGKS) equation [21]. In this preliminary study, Uhlmann’s fidelity is adopted as the scalar performance metric driving the adaptation [31], providing a geometrically meaningful measure of similarity between quantum states. The main contributions of this paper are:
  • the formulation of a continuous-time RCAC-based adaptive controller for LGKS quantum dynamics without requiring a plant model in the control law,
  • the design of an adaptive PID feedback structure driven by fidelity-based error, and
  • numerical validation of the proposed approach for tracking both low-entropy and high-entropy target states in a two-level open quantum system.
@@
The remainder of the paper is organized as follows. Section II describes the LGKS equation in the context of a controlled open quantum system. Section III presents the mathematical foundations of continuous-time RCAC with forgetting factors. Section IV presents numerical simulations validating the proposed adaptive controller on a two-level quantum system. Finally, Section V concludes the paper and outlines directions for future work.
@@ II Problem Formulation
We consider the Lindblad-Gorini-Kossakowski-Sudarshan system as described in [21], which is governed by
${\dot{\rho}}$ $=-i\hbar[H,{{\rho}}]+\sum_{i=1}^{m}\left(L_{i}{{\rho}}L_{i}^{\rm H}-\dfrac{1}{2}\{L_{i}^{\rm H}L_{i},{{\rho}}\}\right),$ (1)
where ${{\rho}}\in\mathbb{C}^{n\times n}$ is the density matrix, the system Hamiltonian $H\in\mathbb{C}^{n\times n}$ is Hermitian, $m$ is the number of damping terms, and for $i\in(1,2,\ldots,m),$ the matrices $L_{i}\in{\mathbb{C}}^{n\times n}$ are the jump operators. Without loss of generality, the units are chosen such that $\hbar=1.$ The diagonal elements of $\rho$}} represent the probability of a quantum event. It is shown in [4, p 123] that, if ${\rm tr}\,{{\rho}}(0)=1,$ then, for all $t>0,$ ${{\rho}}(t)$ given by (1) satisfies ${\rm tr}\,({{\rho}}(t))=1.$ Furthermore, for all $t\geq 0,$ ${{\rho}}(t)$ is positive semidefinite and Hermitian.
In this paper, we consider a two-level quantum dynamic system, that is, $n=2$ and thus ${{\rho}}\in{\mathbb{C}}^{2\times 2}.$ As shown in [2], the Hamilotonian $H$ can be decomposed as
@@
where $H_{0}\in\mathbb{C}^{2\times 2}$ is the free Hamiltonian, $H_{1}\in\mathbb{C}^{2\times 2}$ is the control Hamiltonian, and $u\in\mathbb{R}$ is the scalar control input. The restriction $u\in\mathbb{R}$ ensures that $H$ remains Hermitian at all times.
As shown in Figure 1, the objective is to design an adaptive feedback control law that drives the density matrix $\rho$ to a desired target state $\rho_{\rm d}$ without requiring prior knowledge of $H_{0}$, $H_{1}$, or $L_{i}$ in the control law. We assume that an estimate of the density matrix $\rho$ is available for feedback. In practice, the quantum state is not directly measurable in real time; rather, it must be reconstructed from measurement records using quantum filtering or observer-based techniques, such as those described in [qamar2019observer].
Remark II.1.
Measurement and Observer Assumptions. The present formulation assumes access to an estimate of density $\rho$ for feedback and neglects the effect of measurement on the system dynamics. This corresponds to an ensemble-averaged or mean-field description of the open quantum system and serves as a first step toward evaluating model-free adaptive control at the level of LGKS dynamics. In general, a strict separation principle between quantum state estimation and feedback control does not hold for stochastic quantum systems with measurement backaction. Incorporating stochastic master equations and measurement-induced backaction into the feedback loop is an important direction for future work.
$F(\rho,\rho_{\rm{d}})$ Adaptive Controller $\mathcal{L}\mathcal{G}\mathcal{K}\mathcal{S}$ Homodyne Detection State Observer $\rho_{\rm{}_{d}}$ $e$ $u$ $\rho$ $\rho$
Figure 1: Adaptive feedback control to track desired density matrix $\rho_{\rm d}.$ The estimator, shown in red, consisting of a homodyne detection and a state observer, provides the density matrix $\rho$ for feedback.
@@ II-A Density Error
As shown in [15], the Uhlmann–Jozsa fidelity between two density matrices $\rho_{1},\rho_{2}$ is defined as
@@
It follows from Proposition A.1 that, if $\rho_{1},\rho_{2}$ are the states of (1), then $F(\rho_{1},\rho_{2})\in[0,1].$ Furthermore, $F(\rho_{1},\rho_{2})=1$ if and only if $\rho_{1}=\rho_{2}.$ Finally, as shown in [16], in the case where $\rho_{1},\rho_{2}\in\mathbb{C}^{2\times 2},$ the Uhlmann–Jozsa fidelity is given by
@@
In this paper, we define the scalar density-tracking error as
@@
where $\rho_{\rm{d}}\in\mathbb{C}^{2\times 2}$ is the desired density. This formulation of error ensures that the feedback control law is driven by a scalar variable, thereby avoiding any geometric inconsistencies in the simulation of the LGKS system. Note that the error $e=0$ when $\rho$ is equal to $\rho_{\rm d}.$
@@ II-B Adaptive PID Control
In this work, we consider an adaptive proportional-integral-derivative (PID) controller, which can be written as
@@
where $k_{\rm p}(t),k_{\rm i}(t),$ and $k_{\rm d}(t)$ are the adaptive proportional, integral, and derivative gains. Note that the controller can be written in the regressor form as
@@
$\Phi(t)\stackrel{{\scriptstyle\triangle}}{{=}}\begin{bmatrix}e(t)&\gamma(t)&\dot{e}(t)\end{bmatrix},\quad\theta(t)\stackrel{{\scriptstyle\triangle}}{{=}}\begin{bmatrix}k_{\rm p}(t)\\ k_{\rm i}(t)\\ k_{\rm d}(t)\end{bmatrix}$, (8)
and the integral state $\gamma(t)$ satisfies $\dot{\gamma}(t)=e(t).$
The objective of the adaptive controller is to update the gains in $\theta(t)$ online so that the density error $e(t)$ is driven toward zero. The gain update law is obtained using the continuous-time retrospective cost adaptive control (RCAC) algorithm described in Section III. The PID structure is adopted due to its low-order structure, interpretability, and compatibility with linear parameterizations required by RCAC. More general controller parameterizations, such as full-state feedback, can be incorporated within the same adaptive framework when additional measurements or state features are available.
@@ III Continuous Time Retrospective Cost Adaptive Control
where $x(t)$ is the state, $u(t)$ is the input, $y(t)$ is the measured output, and the vector functions $f$ and $g$ are the dynamics and the output maps. Define the performance variable
@@
Remark III.1.
Specialization to LGKS fidelity tracking. In the present application, the system state $x(t)$ corresponds to the density matrix $\rho(t)$ governed by the LGKS dynamics (1) with controlled Hamiltonian (2). The measured output is chosen as a scalar function of the density matrix, namely the fidelity-based signal $y(t)\stackrel{{\scriptstyle\triangle}}{{=}}F(\rho(t),\rho_{\rm d})$, and the reference is $r(t)\stackrel{{\scriptstyle\triangle}}{{=}}1$. Consequently, the performance variable (11) becomes $z(t)=y(t)-r(t)=F(\rho(t),\rho_{\rm d})-1=-e(t)$, where $e(t)$ is the density error defined in (5). This choice yields a scalar performance variable that is compatible with the linearly parameterized controller (12).
@@
where the regressor matrix $\Phi(t)$ contains the measured data and the vector $\theta(t)$ contains the controller gains to be optimized. Various linear parameterizations of MIMO controllers are described in [goel_2020_sparse_para].
@@
$\Phi_{\rm f}(t)$ $\stackrel{{\scriptstyle\triangle}}{{=}}G_{\rm f}(s)\left[\Phi(t)\right],$ (14)
$u_{\rm f}(t)$ $\stackrel{{\scriptstyle\triangle}}{{=}}G_{\rm f}(s)\left[u(t)\right],$ (15)
@@
Remark III.2.
Filter states and notation. The states $x_{\phi}$ and $x_{u}$ appearing in the filter realizations below are internal states of the dynamic filter $G_{\rm f}(s)$ driven by $\Phi(t)$ and $u(t)$, respectively. These filter states are auxiliary variables used only to compute $\Phi_{\rm f}(t)$ and $u_{\rm f}(t)$ and are not related to the LGKS state $\rho(t)$.
@@
[…]
$\quad+(\Phi(\tau)\hat{\theta})^{\rm T}R_{u}(\Phi(\tau)\hat{\theta})\Big){\rm d}\tau+e^{-\lambda t}\hat{\theta}^{T}R_{\theta}\hat{\theta}$, (20)
where $R_{z},R_{u},$ and $R_{\theta}$ are positive definite weighting matrices of appropriate dimensions, and $\lambda>0$ is the exponential forgetting factor.
The retrospective cost (20) defines an online performance objective for updating the parameter vector $\theta(t)$ based on measured data. We emphasize that minimizing (20) does not, in general, imply closed-loop stability or guarantee that $z(t)\to 0$ for nonlinear systems. In particular, for the LGKS dynamics considered in this paper, the tracking behavior reported in Section IV is empirical and depends on the system, the controller parameterization, and the choice of hyperparameters.
@@
[…]
@@
Next, define, for all $t\geq 0,$ $P(t)\stackrel{{\scriptstyle\triangle}}{{=}}A(t)^{-1}.$ Using the fact that $P(t)A(t)=I,$ it follows that
@@
For the adaptive PID structure in Section II, the regressor $\Phi(t)$ and parameter vector $\theta(t)$ in (12) correspond to the fidelity-based features and gains defined in (6). Thus, the update laws (22) provide an online method to adjust $(k_{\rm p}(t),k_{\rm i}(t),k_{\rm d}(t))$ using only the measured fidelity-based error signal and its integral/derivative features, without requiring explicit knowledge of $H_{0}$, $H_{1}$, or $L_{i}$ in the control law.